OpenStudy (anonymous):

(f+g)(x) =x+3+x-3 =is it 2x ??? or x^2 ????

OpenStudy (mimi_x3):

2x lols

OpenStudy (lalaly):

2x

OpenStudy (angela210793):

x+x=2x x*x=x^2

OpenStudy (anonymous):

actually . were having difficult times regarding this matter . :) thanks!